OpenStudy (anonymous):

can some one solve it f(x) = arctanx/x....f`(0) = ? can some one solve it f(x) = arctanx/x....f`(0) = ? @Mathematics

OpenStudy (anonymous):

so arctanx/x is basically tan inverse/x

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so do it that way

OpenStudy (anonymous):

i have got 0...but im not sure if its correct or not

OpenStudy (anonymous):

whats your value for x that you plugged in?

OpenStudy (anonymous):

f`(0) = ?

OpenStudy (anonymous):

help me with this, how do i differentiate sin(\[\cos(\pi t/4)\]

OpenStudy (anonymous):

yes that is the right answer for you

OpenStudy (anonymous):

(sin(cos(pi*t/4))` = cos ( cos(pi*t/4)*(-sin(pi*t/4)

OpenStudy (anonymous):

differentiate sin(cos(πt/4), let u=cos(πt/4) Dsinu=(cosu)D(cos(πt/4) =(cosu)(-sin(πt/4)(pi/4)Dt

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

=(cos(cos(πt/4) ))(-sin(πt/4)(pi/4)Dt

OpenStudy (anonymous):

i got the question wrong, its actually only \[\cos (pit/4) oops\]